### v1

# Renormalization of the Quantum Stress Tensor Fluctuations and the Limits of Semiclassical Gravity

Abstract
We analyze the expectation value of the energy–momentum tensor and its fluctuations in quantum field theory on curved spacetimes $\langle T_{ab}\rangle$. A necessary condition for the conceptual consistency of semiclassical gravity, where $\langle T_{ab}\rangle$ represent the sources of the Einstein equations, is that the fluctuations of the energy momentum tensor remain small compared to its expectation value. We study the renormalization of both the energy–momentum tensor $\langle T_{ab}(x)\rangle_{\rm ren}$ and the fluctuation tensor $\langle T_{ab}(x)T_{cd}(x)\rangle_{\rm ren}$ for suitable Hadamard states, using the operator product expansion for a free scalar field on a fixed curved background. We show that squeezed vacua—arising naturally in black hole evaporation and in inflationary cosmology—fail to satisfy the semiclassicality criterion.
pacs:
98.80.Es, 04.50.Kd, 03.65.Ta
In the absence of a consistent and well-defined theory of quantum gravity, semiclassical gravity—defined by Einstein’s equations coupled to the expectation value of the energy–momentum tensor of matter in a quantum state—is expected to provide useful insights into phenomena at the interface between gravity and quantum physics. Even though it is not entirely clear whether the dynamical equations of semiclassical gravity can be given a fully consistent interpretation—owing to questions about well-posedness in generic situations stemming from their higher-derivative character Flanagan:1996gw ; Juarez-Aubry:2022qdp —the semiclassical framework nonetheless plays a central role in conceptual discussions. These range from the physics of black-hole evaporation via Hawking radiation, to analyses of the generalized second law, the fate of information, and various issues related to the emergence of structure in cosmology.
In order to explore the regime of applicability of semiclassical gravity, it is necessary to evaluate the quantum uncertainties of the energy–momentum tensor in comparison to the expectation value of that same quantity. The issue would be straightforward were it not for the fact that both the expectation value and the fluctuations are divergent notions that require suitable renormalization in order to yield finite and meaningful results.
The renormalization of expectation values of operators that are quadratic in the fundamental fields and their derivatives (such as the energy–momentum tensor of free fields) is customarily based on the Hadamard subtraction. An axiomatic analysis proposed by Wald Wald:1995yp assures that the recipe is unique up to the higher order curvature terms which (aside from a finite number of them) are naturally expected to be suppressed by appropriate powers of the Planck scale.
For operators built from higher powers of the fundamental fields and their derivatives (as in the case of the tensor product of the energy–momentum tensor with itself), subtracting a state-independent divergent quantity is no longer sufficient to eliminate ultraviolet divergences. In such cases, the more general framework of the operator product expansion (OPE) appears as the consistent path to renormalization WilsonZimmermann1972 . In Hollands:2023txn its generalization to arbitrary (globally hyperbolic) curved spacetimes is proposed and provides a systematic and general procedure which might even be taken as a more fundamental definition of the quantum field theory itself.
The evaluation of the (renormalized) fluctuations of the energy–momentum tensor, and their relevance for assessing the validity of semiclassical gravity, has been considered in the past, beginning with the work of Ford FORD1982238 and later by Ford and Kuo Kuo:1993if , where the problem was studied for quantum fields on Minkowski spacetime. Explicit calculations for certain curved spacetimes admitting a Euclidean continuation were subsequently carried out by Phillips and Hu Phillips:1996em as well as in cosmological settings (see Hu:2008rga and references therein). In these works, several perspectives and definitions of energy–momentum fluctuations were proposed, often invoking non-covariant expressions formulated in terms of specific mean values of components of the energy–momentum tensor, such as the “energy density.” In the early debates, much of the discussion centered on the choice of appropriate criteria for the validity of semiclassical gravity, motivated by the requirement that Minkowski spacetime endowed with the Poincaré-invariant vacuum should constitute a consistent solution of the corresponding field equations.
While our results agree in general lines with these earlier works and in the specific settings they considered, our approach offers a systematic and fully covariant framework that is, in principle, applicable to general spacetimes and general (Hadamard) states. Moreover, the OPE perspective ensures automatically that flat spacetime with matter fields in the Minkowski vacuum state is a consistent solution of semiclassical gravity (see (19)).
The results presented here should be extendable to general quantum field theories. For simplicity, we will focus our discussion on the theory of a massive free scalar field described by the action
$S=\frac{1}{2}\int\sqrt{-g}d^{4}x[\nabla^{a}\phi\nabla_{a}\phi-m^{2}\phi^{2}].$ (1)
As fields are operator-valued distributions in quantum field theory, the product of fields at the same point is ill-defined. The heuristic idea behind the renormalization of quadratic products begins by introducing a point-splitting regularization, in which one considers $\phi(x)\phi(x^{\prime})$ for nearby spacetime events $x$ and $x^{\prime}$, and then restricts attention to states that share a universal UV behavior (the so-called Hadamard states) in the coincidence limit. In this way one can separate the UV and IR pieces of the product as
${\phi(x)\phi(x^{\prime})}\approx C(x,x^{\prime})_{\rm uv}\bf{1}+{(\phi(x)\phi(x^{\prime}))_{\rm ren}},$ (2)
where $\approx$ denotes equality in the limit $x\to x^{\prime}$ when the expression is used to compute expectation values in Hadamard states. Here, $C(x,x^{\prime})_{\rm uv}$ captures the universal divergent behaviour in the coincidence limit, and $(\phi(x)\phi(x^{\prime}))_{\rm ren}$ finite contributions.
The previous heuristic discussion about the separation between UV divergent contributions and finite (physically relevant) long distance trends is at the conceptual heart of the rigorous Hadamard subtraction procedure usually employed to compute the renormalized expectation value of quadratic operators, which in the case of the energy momentum tensor of a free field leads to the expression
$\langle\psi|T_{ab}(x))|\psi\rangle_{\rm ren}=\lim_{x\to x^{\prime}}{{\mathfs{T}}}_{ab^{\prime}}[\langle\psi|\phi(x)\phi(x^{\prime})|\psi\rangle-H(x,x^{\prime})]{}$
where the differential operator ${{\mathfs{T}}}_{ab^{\prime}}$ acting at the tangent planes of the 2-points is given by
${{\mathfs{T}}}_{ab^{\prime}}$ $=$ $\nabla_{a}\nabla_{b^{\prime}}-\frac{1}{2}g_{ab}g^{cd^{\prime}}\nabla_{c}\nabla_{d^{\prime}}-\frac{1}{2}g_{ab}m^{2}$ (3)
where $g_{a}^{b^{\prime}}$ represents the parallel transport operators along the unique geodesic connecting the points $x$ and $x^{\prime}$ within a common convex neighbourhood, and $H(x,x^{\prime})$—which in equation (2) corresponds to $C(x,x^{\prime})_{\rm uv}$—is the Hadamard bi-distributional kernel, constructed from geometrically defined quantities, which is state-independent and captures the UV behavior of the (Hadamard) states considered physically meaningful.
The procedure admits the addition of a state independent divergence free tensor Wald:1995yp whose relevant contributions correspond to a finite renormalization of the cosmological constant and the Newton constant, in addition to two other contributions that could be obtained from variations of an action containing $R^{2}$ and $R_{ab}R^{ab}$ terms. Higher curvature contributions will be suppressed by a mass scale (naturally Planck’s).
In defining the fluctuations of the energy momentum tensor in a given quantum state of the scalar field, the strategy will be to start with the 4-point function $\langle\psi|\phi(x_{1})\phi(x_{2})\phi(x_{3})\phi(x_{4})|\psi\rangle$, from which we can define multi-tensorial quantity
${\mathfs{T}}_{a_{1}b_{2}}{\mathfs{T}}_{c_{3},d_{4}}\langle\psi|\phi(x_{1})\phi(x_{2})\phi(x_{3})\phi(x_{4})|\psi\rangle.$ (4)
The divergence structure of the 4-point function can be illustrated by considering the square of (2), namely
$(\phi(x)\phi(x^{\prime}))^{2}=C^{2}(x,x^{\prime})_{\rm uv}{\bf 1}+$ (5)
$+2C(x,x^{\prime})_{\rm uv}(\phi(x)\phi(x^{\prime}))_{\rm ren}+((\phi(x)\phi(x^{\prime}))^{2}_{\rm ren},{}$
where one can identify a universal UV singular behavior in the first term, a second divergent term which is state dependent (due to the interference with the long distance trend Ford:2000vm ; Ford2000StressTensor ), and the last one which is finite, and physically relevant for the renormalization of (4). The OPE is a systematic approach to extract such finite component which (in the coincidence limit) is defined as $\braket{\psi|T_{ab}(x)T_{cd}(x)|\psi}_{\rm ren}$.
One might be concerned about the ordering issues in taking the coincident limit and other ambiguities. Fortunately, the OPE axiomatic framework ensures that the finite term is uniquely defined up to geometric (state independent) ambiguities similar to those occurring in the renormalization of $T_{ab}(x)$.
The OPE is very general and allows for the renormalization of the expectation value of arbitrary products of fields and their derivatives at a same point. However even when the framework ensures the success of the renormalization program, the specific calculations of quantities of interest, it can become difficult in concrete practical applications.
Fortunately, in case of free fields on curved backgrounds, there is an alternative equivalent avenue for computing the renormalized fluctuations of the energy momentum tensor. With the OPE philosophy in mind, we will find a shortcut using the Wick expansion of the 4-point function associated with a specific choice of vacuum state. This will facilitate explicit calculations on the one hand, and the interpretation of the result in situations where such vacua capture a concrete physical situation.
In order to facilitate the interpretation of the results in situations where a preferred notion of vacuum state is available or physically motivated (as in the case of standard states such as the Hartle–Hawking state, the Unruh state, the in-state for static black hole backgrounds, or the adiabatic vacuum in cosmological settings), it is useful to relate the OPE formalism to the regularization procedure associated with normal ordering. Even though the notion of normal ordering depends on the chosen vacuum state and is therefore highly ambiguous, the quantity $\braket{\psi|T_{ab}(x)T_{cd}(x)|\psi}_{\rm ren}$ will be independent of the normal ordering chosen.
As shown in the appendix A, the 4-point fucntion of the scalar field can be written as
$\braket{\psi|\phi(x_{1})\phi(x_{2})\phi(x_{3})\phi(x_{4})|\psi}=$ (6)
$=\braket{\psi|\phi(x_{1})\phi(x_{2})\phi(x_{3})\phi(x_{4})|\psi}_{\rm ren}+{}$
$+\braket{\psi|\phi(x_{1})\phi(x_{2})|\psi}_{\rm ren}H(x_{3},x_{4})+{\rm other\ pairings}{}$
$+H(x_{1},x_{2})H(x_{3},x_{4})+{\rm other\ pairings}.{}$
where
$\braket{0|\phi(x_{1})\phi(x_{2})|0}_{\rm ren}\equiv\braket{0|\phi(x_{1})\phi(x_{2})|0}-H(x_{1},x_{2}).$ (7)
and the finite renormalized component of the 4-point function is given by
$\!\!\!\!\!\!\!\!\!\braket{\psi|\phi(x_{1})\phi(x_{2})\phi(x_{3})\phi(x_{4})|\psi}_{\rm ren}\equiv$ (8)
$\!\!\!\!\!\!\!\!\!\equiv\braket{\psi|\!:\!\phi(x_{1})\phi(x_{2})\phi(x_{3})\phi(x_{4})\!:\!|\psi}+{}$
$\!\!\!\!\!\!\!\!\!\braket{\psi|\!:\!\phi(x_{1})\phi(x_{2})\!:\!|\psi}\braket{0|\phi(x_{3})\phi(x_{4})|0}_{\rm ren}\!+{\rm other\ pairings}{}$
$\!\!\!\!\!\!\!\!\!\braket{0|\phi(x_{1})\phi(x_{2})|0}_{\rm ren}\braket{0|\phi(x_{3})\phi(x_{4})|0}_{\rm ren}+{\rm other\ pairings}.{}$
Performing the multiple Taylor expansion of the renormalized quantities (those with the subindex “ren”) in (6) (for example writing $\braket{\psi|\phi(x)\phi(x+y)|\psi}_{\rm ren}=\braket{\psi|\phi(x)^{2}|\psi}_{\rm ren}+y^{\mu}\partial_{\mu}\braket{\psi|\phi(x)^{2}|\psi}_{\rm ren}+\cdots$) we would obtain the OPE for the 4-point function of a free field on a curved background. As each of the divergent terms in (6) are canonical, the finite term (8) does not depend on the choice of the vacuum state used in the normal ordering. The renormalized square of the energy momentum tensor can now be given a meaning as
$\braket{\psi|T_{ab}(x)T_{cd}(x)|\psi}_{\rm ren}{}$
$=\!\!\!\!\!\!\!\lim_{x_{1},x_{2},x_{3},x_{4}\to x}\!\!\!\!\!\!\!{\mathfs{T}}_{a_{1}b_{2}}{\mathfs{T}}_{c_{3}d_{4}}\braket{\psi|\phi(x_{1})\phi(x_{2})\phi(x_{3})\phi(x_{4})|\psi}_{\rm ren}{}$
We can simplify the discussion of the fluctuations of the energy momentum tensor by focusing on an extreme idealization of a massive scalar field in its large mass limit. In that case
$\braket{T_{ab}}_{\rm ren}\approx m^{2}\braket{\phi^{2}}_{\rm ren}g_{ab}.$ (9)
Focusing on the trace, (8) implies that for any squeezed state (defined as a Bogoliubov transformation of an arbitrary vacuum state) such that $\braket{\phi^{2}}_{\rm ren}\not=0$, we have
${\Delta T^{2}_{\rm ren}}\equiv{\braket{T^{2}}_{\rm ren}-\braket{T}^{2}_{\rm ren}}\approx 2{\braket{T}_{\rm ren}^{2}}.$ (10)
In what follows, we will show that the same result holds in full generality within a qualitatively realistic, yet simplified, 2-dimensional model.
I Explicit calculation in 2d
The renormalized expectation value of the energy momentum tensor of a conformally invariant scalar field in 2d backgrounds can be written as
$\braket{\psi|T_{\pm\pm}(x)|\psi}_{\rm ren}=\lim_{x_{1},x_{2}\to x}\braket{\psi|\partial_{x}^{\pm}\phi(x_{1})\partial_{x}^{\pm}\phi(x_{2})|\psi}_{\rm ren}{}$
$\equiv\lim_{x_{1},x_{2}\to x}\frac{\partial X^{\pm}_{1}}{\partial x^{\pm}_{1}}\frac{\partial X^{\pm}_{2}}{\partial x^{\pm}_{2}}\braket{\psi|:\partial_{X}^{\pm}\phi(x_{1})\partial_{X}^{\pm}\phi(x_{2}):_{X}|\psi}{},$
[…]
$\braket{\psi|T_{\pm\pm}(x)|\psi}_{\rm ren}=\braket{\psi|:T_{\pm\pm}:_{x}|\psi}+\frac{\hbar}{12\pi}\tau_{\pm}(\Omega),$ (11)
[…]
$\tau_{\pm}(\Omega)\equiv\left(\partial^{2}_{\pm}\log(\Omega)-\partial_{\pm}\log(\Omega)\partial_{\pm}\log(\Omega)\right),$ (12)
with $\Omega$ the conformal factor appearing in the metric
$ds^{2}=-\Omega^{2}dx^{+}dx^{-}.$ (13)
One can show that the left hand side of (11) transforms covariantly under diffeomorphisms (with each term on the right transforming with the Schwarzian derivative with opposite signs). The conformal anomaly (see Appendix) fixes
$\braket{\psi|T_{+-}|\psi}=-\frac{\hbar}{12\pi}\partial_{+}\partial_{-}\log(\Omega).$ (14)
For the fluctuations of $T_{ab}$ we proceed along similar lines, defining their renormalized value as the normal-ordered value with respect to the local freely falling vacuum, namely
$\braket{\psi|T_{\pm\pm}(x)T_{\pm\pm}(x)|\psi}_{\rm ren}=$ (15)
$\equiv\lim_{x_{1},x_{2},x_{3},x_{4}\to x}\frac{\partial X^{\pm}_{1}}{\partial x^{\pm}_{1}}\frac{\partial X^{\pm}_{2}}{\partial x^{\pm}_{2}}\frac{\partial X^{\pm}_{3}}{\partial x^{\pm}_{3}}\frac{\partial X^{\pm}_{4}}{\partial x^{\pm}_{4}}\times{}$
$\times\braket{\psi|:\partial_{X}^{\pm}\phi(x_{1})\partial_{X}^{\pm}\phi(x_{2})\partial_{X}^{\pm}\phi(x_{3})\partial_{X}^{\pm}\phi(x_{4}):_{X}|\psi}{},$
where, as in (I), we denote by $_{X}$ the normal ordered prescription based on the vacuum notion constructed out of the positive frequency notion associated with freely falling coordinates null coordinates $X^{\pm}$. In Appendix B we explicitly show that the previous procedure is equivalent to the OPE perspective advocated in Hollands:2023txn . We also show that the relevant renormalized fluctuations are given by
$\braket{\psi|T_{\pm\pm}T_{\pm\pm}|\psi}_{\rm ren}\equiv\braket{\psi|:T_{\pm\pm}T_{\pm\pm}:|\psi}$ (16)
$+\frac{\hbar}{2\pi}\braket{\psi|:T_{\pm\pm}:|\psi}\tau_{\pm}(\Omega)+\frac{\hbar^{2}}{48\pi^{2}}\tau_{\pm}^{2}(\Omega),{}$
whose covariance is granted by construction from the definition (15). The same method allows to calculate the $n$-point functions of $\partial_{\pm}\phi$ in a closed manner.
We can now apply the previous formulae in order to compute the renormalized fluctuations of the energy momentum tensor. We define
[…]
$\frac{\hbar}{12\pi}\tau_{\pm}<T_{\pm\pm}(\phi_{\rm cl})$ (17)
[…]
which implies that such vacua cannot be considered semiclassical states.
The 2d simplified setting, while qualitatively realistic, facilitates also the definition of a coordinate independent semiclassicality criterion such as (17). In four dimensions, however, covariance would require a more careful treatment. One could for instance demand that for all sets of orthonormal tetrads $\{e^{a}_{I}\}$ the quantities $T_{IJ}\equiv e^{a}_{I}e^{b}_{J}\hat{T}_{ab}$ would have quantum uncertainties that are small compared with the corresponding expectation values. A less restrictive criteria would demand only that certain scalar objects such as say the trace $T=g^{ab}\hat{T}_{ab}$ (as in (10)) be such that its expectation value dominates over the corresponding uncertainties.
In view of the previous discussion we propose as a precise criterion for the validity of semiclassical gravity at the spacetime point $x$ is that
$\left|\braket{T_{IJ}^{2}(x)}_{\rm ren}\right|\leq\epsilon\braket{T_{IJ}(x)}_{\rm ren}^{2}$ (19)
with $\epsilon<1$ for any tetrad. Note that Minkowski with the Poincare invariant vacuum satisfies this at every point.
II Discussion
In the context of 2d black holes, states for which (18) applies include emblematic cases such as the Boulware vacuum, the Hartle-Hawking state, the Unruh state, as well as the in-state (representing the particle creation in the setting of gravitational collapse). This could have been naively anticipated from the Gaussian nature of such states as a general consequence of the Wick theorem. Here we have shown that the expectation remains correct after renormalization. We note that in 4d the a field $\Phi$ in the spherically symmetric modes satisfies the Klein-Gordon equation (for a massive or massless cases) at the black hole horizon if $\Phi=\phi/r$ with $\phi$ satisfying the conformally invariant 2d equation of the field in the previous section. Moreover, the spherical mode contribution to the energy momentum tensor at the black hole horizon, for the relevant component encoding the radiation flux across the horizon, is $T^{\rm sph}_{vv}=T^{\rm 2d}_{vv}/r^{2}$ (see Appendix C). Even when the in and Unruh quantum states would change in the generalization to 4d due to the backscattering by the effective potential that is non trivial outside of the horizon, one expects that the previous identification would be qualitatively sensible for the expectation values and fluctuations of the above quantity. This fact, combined with our discussion of the 4-point function OPE in 4d, indicate that the conclusions should remain valid in any dimension (even when explicit evaluation might be difficult).
[…]
Even when weak, their quantum gravitational effects can be significant in situations of interest, such as in the vicinity of macroscopic black holes. We show that the fluctuations of the energy–momentum tensor are of the same order of magnitude as the expectation values that drive black hole evaporation. This raises important questions about the validity of the semiclassical heuristic picture in analyzing the long-term evolution of such systems. In particular, our analysis seems to align with the insightful physical argument of Page Page:1979tc , which implies large uncertainties in the black hole’s position during evaporation, a conclusion strongly at odds with the semiclassical description.
The idea that semiclassical gravity cannot be applied in general situations is usually associated with scenarios involving macroscopic superpositions of gravitational sources at different locations (an idea that could, in principle, admit experimental confirmation Page:1981aj ). Here we show that, after proper renormalization, large fluctuations in gravitational sources are expected to be present in a broad class of quantum states of matter, including familiar and emblematic physical idealizations. This raises important questions regarding the status of many conceptual claims about the nature of quantum effects in the contexts of black-hole physics and primordial cosmology. It also reinforces the view that some of these issues can be properly addressed only within a framework in which the quantum nature of the gravitational degrees of freedom is fully taken into account.
III Acknowledgments
We are grateful to B. Kay, B. Juarez Aubry for clarifying discussions in early stages of this work, and to M. Knecht and I. Agullo for helpful exchanges. DS acknowledges support from the Proyect PAPIIT No IG100124, and a sabbatical year Fellowship from UNAM, through a PASPA/DGAPA and the hospitality of the Centre de Physique Théorique in Marseille, and the Departament de Física Quantica i Astrofísica Universitat de Barcelona.
Appendix A Four point function in 4d
[…]
$\braket{\psi|\phi(x_{1})\phi(x_{2})\phi(x_{3})\phi(x_{4})|\psi}=\braket{\psi|\!:\!\phi(x_{1})\phi(x_{2})\phi(x_{3})\phi(x_{4})\!:\!|\psi}+{}$
$\braket{\psi|\!:\!\phi(x_{1})\phi(x_{2})\!:\!|\psi}\braket{0|\phi(x_{3})\phi(x_{4})|0}+\braket{\psi|\!:\!\phi(x_{1})\phi(x_{3})\!:\!|\psi}\braket{0|\phi(x_{2})\phi(x_{4})|0}+\braket{\psi|\!:\!\phi(x_{1})\phi(x_{4})\!:\!|\psi}\braket{0|\phi(x_{2})\phi(x_{3})|0}+{}$
$\braket{\psi|\!:\!\phi(x_{2})\phi(x_{3})\!:\!|\psi}\braket{0|\phi(x_{1})\phi(x_{4})|0}+\braket{\psi|\!:\!\phi(x_{2})\phi(x_{4})\!:\!|\psi}\braket{0|\phi(x_{1})\phi(x_{3})|0}+\braket{\psi|\!:\!\phi(x_{3})\phi(x_{4})\!:\!|\psi}\braket{0|\phi(x_{1})\phi(x_{2})|0}+{}{}$
$\braket{0|\phi(x_{2})\phi(x_{3})|0}\braket{0|\phi(x_{1})\phi(x_{4})|0}+\braket{0|\phi(x_{2})\phi(x_{4})|0}\braket{0|\phi(x_{1})\phi(x_{3})|0}+\braket{0|\phi(x_{3})\phi(x_{4})|0}\braket{0|\phi(x_{1})\phi(x_{2})|0}.{}$
[…]
$\braket{0|\phi(x_{1})\phi(x_{2})|0}_{\rm ren}\equiv\braket{0|\phi(x_{1})\phi(x_{2})|0}-H(x_{1},x_{2}).$ (20)
Using the previous equation in order to replace the vacuum expectation values in terms of renormalized 2-point function and the Hadamard bi distribution we get
$\braket{\psi|\phi(x_{1})\phi(x_{2})\phi(x_{3})\phi(x_{4})|\psi}=\braket{\psi|\!:\!\phi(x_{1})\phi(x_{2})\phi(x_{3})\phi(x_{4})\!:\!|\psi}+$ (21)
$+\braket{\psi|\!:\!\phi(x_{1})\phi(x_{2})\!:\!|\psi}\braket{0|\phi(x_{3})\phi(x_{4})|0}_{\rm ren}+\braket{\psi|\!:\!\phi(x_{1})\phi(x_{3})\!:\!|\psi}\braket{0|\phi(x_{2})\phi(x_{4})|0}_{\rm ren}+{}$
$+\braket{\psi|\!:\!\phi(x_{1})\phi(x_{4})\!:\!|\psi}\braket{0|\phi(x_{2})\phi(x_{3})|0}_{\rm ren}+\braket{\psi|\!:\!\phi(x_{2})\phi(x_{3})\!:\!|\psi}\braket{0|\phi(x_{1})\phi(x_{4})|0}_{\rm ren}+{}$
$+\braket{\psi|\!:\!\phi(x_{2})\phi(x_{4})\!:\!|\psi}\braket{0|\phi(x_{1})\phi(x_{3})|0}_{\rm ren}+\braket{\psi|\!:\!\phi(x_{3})\phi(x_{4})\!:\!|\psi}\braket{0|\phi(x_{1})\phi(x_{2})|0}_{\rm ren}+{}{}$
$+\braket{0|\phi(x_{2})\phi(x_{3})|0}_{\rm ren}\!\braket{0|\phi(x_{1})\phi(x_{4})|0}_{\rm ren}+\braket{0|\phi(x_{2})\phi(x_{4})|0}_{\rm ren}\!\braket{0|\phi(x_{1})\phi(x_{3})|0}_{\rm ren}+{}$
$+\braket{0|\phi(x_{3})\phi(x_{4})|0}_{\rm ren}\!\braket{0|\phi(x_{1})\phi(x_{2})|0}_{\rm ren}+{}$
$+\braket{\psi|\!:\!\phi(x_{1})\phi(x_{2})\!:\!|\psi}H(x_{3},x_{4})+\braket{\psi|\!:\!\phi(x_{1})\phi(x_{3})\!:\!|\psi}H(x_{2},x_{4})+\braket{\psi|\!:\!\phi(x_{1})\phi(x_{4})\!:\!|\psi}H(x_{2},x_{3})+{}$
$+\braket{\psi|\!:\!\phi(x_{2})\phi(x_{3})\!:\!|\psi}H(x_{1},x_{4})+\braket{\psi|\!:\!\phi(x_{2})\phi(x_{4})\!:\!|\psi}H(x_{1},x_{3})+\braket{\psi|\!:\!\phi(x_{3})\phi(x_{4})\!:\!|\psi}H(x_{1},x_{2})+{}{}$
$+H(x_{2},x_{3})\braket{0|\phi(x_{1})\phi(x_{4})|0}_{\rm ren}+H(x_{2},x_{4})\braket{0|\phi(x_{1})\phi(x_{3})|0}_{\rm ren}+H(x_{3},x_{4})\braket{0|\phi(x_{1})\phi(x_{2})|0}_{\rm ren}{}$
$+H(x_{1},x_{2})\braket{0|\phi(x_{3})\phi(x_{4})|0}_{\rm ren}+H(x_{1},x_{3})\braket{0|\phi(x_{2})\phi(x_{4})|0}_{\rm ren}+H(x_{1},x_{4})\braket{0|\phi(x_{2})\phi(x_{3})|0}_{\rm ren}+{}$
$+H(x_{1},x_{2})H(x_{3},x_{4})+H(x_{1},x_{3})H(x_{2},x_{4})+H(x_{1},x_{4})H(x_{2},x_{3}).{}$
Inspection of the previous equation allows for identifying the term that is free of UV divergences, namely
$\braket{\psi|\phi(x_{1})\phi(x_{2})\phi(x_{3})\phi(x_{4})|\psi}_{\rm ren}\equiv\braket{\psi|\!:\!\phi(x_{1})\phi(x_{2})\phi(x_{3})\phi(x_{4})\!:\!|\psi}+$ (22)
$+\braket{\psi|\!:\!\phi(x_{1})\phi(x_{2})\!:\!|\psi}\braket{0|\phi(x_{3})\phi(x_{4})|0}_{\rm ren}+\braket{\psi|\!:\!\phi(x_{1})\phi(x_{3})\!:\!|\psi}\braket{0|\phi(x_{2})\phi(x_{4})|0}_{\rm ren}+{}$
$+\braket{\psi|\!:\!\phi(x_{1})\phi(x_{4})\!:\!|\psi}\braket{0|\phi(x_{2})\phi(x_{3})|0}_{\rm ren}+\braket{\psi|\!:\!\phi(x_{2})\phi(x_{3})\!:\!|\psi}\braket{0|\phi(x_{1})\phi(x_{4})|0}_{\rm ren}+{}$
$+\braket{\psi|\!:\!\phi(x_{2})\phi(x_{4})\!:\!|\psi}\braket{0|\phi(x_{1})\phi(x_{3})|0}_{\rm ren}+\braket{\psi|\!:\!\phi(x_{3})\phi(x_{4})\!:\!|\psi}\braket{0|\phi(x_{1})\phi(x_{2})|0}_{\rm ren}+{}{}$
$+\braket{0|\phi(x_{1})\phi(x_{2})|0}_{\rm ren}\braket{0|\phi(x_{3})\phi(x_{4})|0}_{\rm ren}+\braket{0|\phi(x_{1})\phi(x_{3})|0}_{\rm ren}\braket{0|\phi(x_{2})\phi(x_{4})|0}_{\rm ren}+{}$
$+\braket{0|\phi(x_{1})\phi(x_{4})|0}_{\rm ren}\braket{0|\phi(x_{2})\phi(x_{3})|0}_{\rm ren},{}$
which, when replaced in (21), yields
$\braket{\psi|\phi(x_{1})\phi(x_{2})\phi(x_{3})\phi(x_{4})|\psi}=\braket{\psi|\phi(x_{1})\phi(x_{2})\phi(x_{3})\phi(x_{4})|\psi}_{\rm ren}+{}$
$+(\braket{\psi|\!:\!\phi(x_{1})\phi(x_{2})\!:\!|\psi}+\braket{0|\phi(x_{1})\phi(x_{2})|0}_{\rm ren})H(x_{3},x_{4})+\left(\braket{\psi|\!:\!\phi(x_{1})\phi(x_{3})\!:\!|\psi}+\braket{0|\phi(x_{1})\phi(x_{3})|0}_{\rm ren}\right)H(x_{2},x_{4})+{}$
$+(\braket{\psi|\!:\!\phi(x_{1})\phi(x_{4})\!:\!|\psi}+\braket{0|\phi(x_{1})\phi(x_{4})|0}_{\rm ren})H(x_{2},x_{3})+(\braket{\psi|\!:\!\phi(x_{2})\phi(x_{3})\!:\!|\psi}+\braket{0|\phi(x_{2})\phi(x_{3})|0}_{\rm ren})H(x_{1},x_{4})+{}$
$+(\braket{\psi|\!:\!\phi(x_{2})\phi(x_{4})\!:\!|\psi}+\braket{0|\phi(x_{2})\phi(x_{4})|0}_{\rm ren})H(x_{1},x_{3})+(\braket{\psi|\!:\!\phi(x_{3})\phi(x_{4})\!:\!|\psi}+\braket{0|\phi(x_{3})\phi(x_{4})|0}_{\rm ren})H(x_{1},x_{2})+{}{}$
$+H(x_{1},x_{2})H(x_{3},x_{4})+H(x_{1},x_{3})H(x_{2},x_{4})+H(x_{1},x_{4})H(x_{2},x_{3}).$ (23)
Despite the explicit appearance of normal-ordered contributions in the renormalized correlation function (22), one can see that it is indeed independent of the choice of the reference vacuum state used to define normal ordering. This becomes clear by performing one final rearrangement of (6) as follows:
$\braket{\psi|\phi(x_{1})\phi(x_{2})\phi(x_{3})\phi(x_{4})|\psi}=\braket{\psi|\phi(x_{1})\phi(x_{2})\phi(x_{3})\phi(x_{4})|\psi}_{\rm ren}+{}$
$+(\braket{\psi|\phi(x_{1})\phi(x_{2})|\psi}-H(x_{1},x_{2}))H(x_{2},x_{3})+\left(\braket{\psi|\phi(x_{1})\phi(x_{3})|\psi}-H(x_{1},x_{3})\right)H(x_{2},x_{4})+{}$
$+(\braket{\psi|\phi(x_{1})\phi(x_{4})|\psi}-H(x_{1},x_{4}))H(x_{2},x_{3})+(\braket{\psi|\phi(x_{2})\phi(x_{3})|\psi}-H(x_{2},x_{3}))H(x_{1},x_{4})+{}$
$+(\braket{\psi|\phi(x_{2})\phi(x_{4})|\psi}-H(x_{2},x_{4}))H(x_{1},x_{3})+(\braket{\psi|\phi(x_{3})\phi(x_{4})|\psi}-H(x_{3},x_{4}))H(x_{1},x_{2})+{}{}$
$+H(x_{1},x_{2})H(x_{3},x_{4})+H(x_{1},x_{3})H(x_{2},x_{4})+H(x_{1},x_{4})H(x_{2},x_{3}).$ (24)
or
$\braket{\psi|\phi(x_{1})\phi(x_{2})\phi(x_{3})\phi(x_{4})|\psi}=\braket{\psi|\phi(x_{1})\phi(x_{2})\phi(x_{3})\phi(x_{4})|\psi}_{\rm ren}+{}$
$\braket{\psi|\phi(x_{1})\phi(x_{2})|\psi}_{\rm ren}H(x_{2},x_{3})+\braket{\psi|\phi(x_{1})\phi(x_{3})|\psi}_{\rm ren}H(x_{2},x_{4})+{}$
$\braket{\psi|\phi(x_{1})\phi(x_{4})|\psi}_{\rm ren}H(x_{2},x_{3})+\braket{\psi|\phi(x_{2})\phi(x_{3})|\psi}_{\rm ren}H(x_{1},x_{4})+{}$
$\braket{\psi|\phi(x_{2})\phi(x_{4})|\psi}_{\rm ren}H(x_{1},x_{3})+\braket{\psi|\phi(x_{3})\phi(x_{4})|\psi}_{\rm ren}H(x_{1},x_{2})+{}{}$
$+H(x_{1},x_{2})H(x_{3},x_{4})+H(x_{1},x_{3})H(x_{2},x_{4})+H(x_{1},x_{4})H(x_{2},x_{3}).$ (25)
The left hand side is obviously independent of the choice of reference vacuum. All the terms on the right hand side, except from the first one, are also manifestly independent of the reference vacuum. Thus, the first term $\braket{\psi|\phi(x_{1})\phi(x_{2})\phi(x_{3})\phi(x_{4})|\psi}_{\rm ren}$ is independent of the reference vacuum state used in the expression (22).
Appendix B Renormalization details in a conformal 2d quantum field theory
The analysis of the previous section could be adapted directly to the computations performed here. However, because the two-point function of a massless scalar field $\phi$ is not well defined (due to infrared divergences) and due to special features associated with the conformal symmetry of the model, it is more convenient to carry out the renormalization of the energy–momentum tensor (and its fluctuations) directly in terms of the well-behaved and unambiguous correlation functions of $d\phi$ instead. Added interest of this section is that it starts with an explicit a self-contained, very concise, and explicit review of the covariant renormalization the expectation value $T_{\mu\nu}$ in 2d (more details of which can be found for instance in Fabbri:2005mw and references therein).
In 2d there is a Fock representation for each conformal coordinate system with the metric of the form (13) and two-point function in the associated vacuum is given by
$\braket{0_{x}|\phi(x_{1})\phi(x_{2})|0_{x}}\approx\frac{\hbar}{4\pi}\left(2\gamma+\log\left[\frac{|(x^{+}_{1}-x^{+}_{2})(x^{-}_{1}-x^{-}_{2})|}{L^{2}}\right]\right),$ (26)
where $L$ is an infrared cutoff, and $\approx$ indicates that the expression is an approximation whose corrections play no role in the formulas that follow—those involving derivatives of the above correlation function—in the coincidence limit. The cut-off dependence disappears when considering the 2-point function of $d\phi$ instead, which is the one of interest in the definition of the energy momentum tensor for the conformal scalar field. Namely,
$\braket{0_{x}|\partial_{\pm}\phi(x_{1})\partial_{\pm}\phi(x_{2})|0_{x}}\approx-\frac{\hbar}{12\pi}\frac{1}{|x^{\pm}_{1}-x^{\pm}_{2}|^{2}}.$ (27)
On can regularize the expectation value of the energy momentum tensor by a point splitting procedure and the corresponding subtraction
$\braket{\psi|:T_{\pm\pm}(x):_{x}|\psi}$ $=$ $\lim_{\Delta\to 0}\left[\braket{\psi|\partial_{\pm}\phi(x)\partial_{\pm}\phi(x+\Delta)|\psi}-\braket{0|\partial_{\pm}\phi(x)\partial_{\pm}\phi(x+\Delta)|0}\right]={}$ (28)
$=$ $\lim_{\Delta\to 0}\left[\braket{\psi|\partial_{\pm}\phi(x)\partial_{\pm}\phi(x+\Delta)|\psi}+\frac{\hbar}{4\pi}\frac{1}{\Delta^{2}}\right],$
[…]
We notice that two normal ordering prescriptions are related by
$\braket{\psi|:T_{\pm\pm}(x):_{x}|\psi}$ - $\left(\frac{\partial y^{\pm}}{\partial x^{\pm}}\right)^{2}\braket{\psi|:T_{\pm\pm}(y):_{y}|\psi}=$
$=$ $\lim_{\Delta\to 0}\frac{\hbar}{4\pi}\left(\frac{1}{|\Delta^{\pm}|^{2}}-\frac{\frac{\partial y^{\pm}}{\partial x^{\pm}}(x)\frac{\partial y^{\pm}}{\partial x^{\pm}}(x+\Delta)}{|y^{\pm}(x)-y^{\pm}(x+\Delta)|^{2}}\right)=-\frac{\hbar}{24\pi}\{y^{\pm},x^{\pm}\},{}$
where
$\{y,x\}\equiv\frac{\frac{d^{3}y}{dx^{3}}}{\frac{dy}{dx}}-\frac{3}{2}\left(\frac{\frac{d^{2}y}{dx^{2}}}{\frac{dy}{dx}}\right)^{2}$ (30)
is the so-called Schwarzian derivative.
B.1 Calculation of $\braket{\psi|T_{\pm\pm}(x)|\psi}_{\rm ren}$
One uses the previous relation between different notions of normal ordered regularizations in order to construct the covariant renormalization prescription (I) which leads to
$\braket{\psi|T_{\pm\pm}(x)|\psi}_{\rm ren}=\braket{\psi|:T_{\pm\pm}:_{x}|\psi}-\frac{\hbar}{12\pi}\left(\partial_{\pm}\log(\Omega)\partial_{\pm}\log(\Omega)-\partial^{2}_{\pm}\log(\Omega)\right),$ (31)
[…]
$\frac{\hbar}{24\pi}\{X^{\pm},x^{\pm}\}=-\frac{\hbar}{12\pi}\left(\partial_{\pm}\log(\Omega)\partial_{\pm}\log(\Omega)-\partial^{2}_{\pm}\log(\Omega)\right)$ (32)
where $X^{\pm}$ are Riemann normal coordinates Fabbri:2005mw
$X^{\pm}_{p}\equiv\Omega(x^{\pm}-x^{\pm}(p))+\Omega\,\partial_{\pm}\log(\Omega)(x^{\pm}-x^{\pm}(p))^{2}+\Omega\left(\frac{1}{3}\partial_{\pm}^{2}(\log(\Omega))+\frac{2}{3}(\partial_{\pm}\log(\Omega))^{2}\right)(x^{\pm}-x^{\pm}(p))^{3}+\cdots$ (33)
with $\Omega$ is the conformal factor in (13). These coordinates are such that the metric (13) takes the form
$ds^{2}=-dX^{+}dX^{-}\left(1+\frac{1}{4}\Omega^{4}R(X^{+}-X^{+}(p))(X^{-}-X^{-}(p))\right)+{\mathfs{O}}((X-X(p))^{3}),$ (34)
in a local neighbourhood of every point $p\in{\mathfs{M}}$. The classical expression of $T_{+-}$ vanishes identically due conformal invariance. The divergence $\nabla^{a}T_{ab}=g^{ca}(\partial_{c}T_{ab}-\Gamma^{d}_{ca}T_{db})$ in the conformal coordinates become $g^{+-}(\partial_{+}T_{-b}+\partial_{-}T_{+b})$ as the only non vanishing Christoffel symbols $\Gamma_{\pm\pm}^{\pm}$ drop due to $T_{+-}=0$. Using (31) we get
$\nabla^{a}\braket{\psi|T_{a\pm}|\psi}_{\rm ren}=2\Omega^{-2}\partial_{\mp}\braket{\psi|T_{\pm\pm}|\psi}=-\frac{\hbar}{12\pi}\partial_{\mp}\left(\partial_{\pm}\log(\Omega)\partial_{\pm}\log(\Omega)-\partial^{2}_{\pm}\log(\Omega)\right)=-\frac{\hbar}{48\pi}\partial_{\pm}R,$ (35)
where we used that Wald:1984rg
$R=-2g^{ab}\nabla_{a}\nabla_{b}\log{\Omega}=8\Omega^{-2}\partial_{+}\partial_{-}\log(\Omega).$ (36)
The previous equation can be written covariantly as
$\nabla^{a}\braket{\psi|T_{ab}|\psi}_{\rm ren}=-\frac{\hbar}{48\pi}\nabla_{b}R,$ (37)
which we call the diffeomorphism anomaly. One can absorb the anomaly defining a modified renormalization prescription via
$\braket{\psi|\tilde{T}_{ab}|\psi}_{\rm ren}\equiv\braket{\psi|T_{ab}|\psi}_{\rm ren}+\frac{\hbar}{48\pi}Rg_{ab}$ (38)
which is divergence free, yet not traceless
$g^{ab}\braket{\psi|\tilde{T}_{ab}|\psi}_{\rm ren}=\frac{\hbar}{24\pi}R,$ (39)
which is the so-called trace anomaly. Both posibilities are also realized in 4d. In the first case diffeomorphisms are broken down to volume-preserving diffeomorphism and the gravity field equations are those of semiclassical trace-free gravity (unimodular gravity). In the second case the gravity field equations are semiclassical Einsteins equations. Both alternatives give the same gravitational dynamics Josset:2016vrq ; Amadei:2021aqd .
B.2 Calculation of $\braket{\psi|T_{\pm\pm}T_{\pm\pm}|\psi}_{\rm ren}$
For the 4-point function needed in the point-splitting-regularization of the square of the energy momentum tensor we shall introduce a simplifying notation where
$\braket{\psi|\partial_{\pm}\phi(x_{1})\partial_{\pm}\phi(x_{2})\partial_{\pm}\phi(x_{3})\partial_{\pm}\phi(x_{4})|\psi}\equiv\braket{1_{\pm}2_{\pm}3_{\pm}4_{\pm}}^{x}_{\psi}{}$
$\braket{\psi|:\partial_{\pm}\phi(x_{1})\partial_{\pm}\phi(x_{2})\partial_{\pm}\phi(x_{3})\partial_{\pm}\phi(x_{4}):|\psi}\equiv\,\braket{:1_{\pm}2_{\pm}3_{\pm}4_{\pm}\!:}_{\psi}^{x},$ (40)
[…]
$\braket{0_{x}|\partial_{\pm}\phi(x_{1})\partial_{\pm}\phi(x_{2})|0_{x}}=-\frac{\hbar}{12\pi}\frac{1}{|x^{\pm}_{1}-x^{\pm}_{2}|^{2}}\equiv\xi_{12}^{x^{\pm}}=\xi_{21}^{x^{\pm}}$ (41)
With this notation, the Wick theorem implies that
$\braket{1_{\pm}2_{\pm}3_{\pm}4_{\pm}}^{x}_{\psi}=\braket{:1_{\pm}2_{\pm}3_{\pm}4_{\pm}\!:}^{x}_{\psi}+{}$
$+\braket{:1_{\pm}2_{\pm}\!:}^{x}_{\psi}\xi_{34}^{x^{\pm}}+\braket{:1_{\pm}3_{\pm}\!:}^{x}_{\psi}\xi_{24}^{x^{\pm}}+\braket{:1_{\pm}4_{\pm}\!:}^{x}_{\psi}\xi_{23}^{x^{\pm}}+\braket{:2_{\pm}3_{\pm}\!:}^{x}_{\psi}\xi_{14}^{x^{\pm}}+\braket{:2_{\pm}4_{\pm}\!:}^{x}_{\psi}\xi_{13}^{x^{\pm}}+\braket{:3_{\pm}4_{\pm}\!:}^{x}_{\psi}\xi_{12}^{x^{\pm}}+{}$
$+\xi_{12}^{x^{\pm}}\xi_{34}^{x^{\pm}}+\xi_{13}^{x^{\pm}}\xi_{24}^{x^{\pm}}+\xi_{14}^{x^{\pm}}\xi_{23}^{x^{\pm}}$ (42)
Using the transformation rule for the 4-point function
$\braket{1_{\pm}2_{\pm}3_{\pm}4_{\pm}}^{x}_{\psi}=\frac{\partial y^{\pm}_{1}}{\partial x^{\pm}_{1}}\frac{\partial y^{\pm}_{2}}{\partial x^{\pm}_{2}}\frac{\partial y^{\pm}_{3}}{\partial x^{\pm}_{3}}\frac{\partial y^{\pm}_{4}}{\partial x^{\pm}_{4}}\braket{1_{\pm}2_{\pm}3_{\pm}4_{\pm}}^{y}$ (43)
Following the same geometric idea used to define the energy momentum tensor, we define the renormalized 4-point correlations function, finite in the coincidence limit, by relating it to the value of the normal ordered 4-point function calculated in the (geometrically defined) local freely falling frame $X^{\pm}$, namely
$\braket{1_{\pm}2_{\pm}3_{\pm}4_{\pm}}^{x}_{\psi{\rm ren}}=\frac{\partial X^{\pm}_{1}}{\partial x^{\pm}_{1}}\frac{\partial X^{\pm}_{2}}{\partial x^{\pm}_{2}}\frac{\partial X^{\pm}_{3}}{\partial x^{\pm}_{3}}\frac{\partial X^{\pm}_{4}}{\partial x^{\pm}_{4}}\braket{:1_{\pm}2_{\pm}3_{\pm}4_{\pm}\!:}^{X}_{\psi}\equiv J_{1}J_{2}J_{3}J_{4}\braket{:1_{\pm}2_{\pm}3_{\pm}4_{\pm}\!:}^{X}_{\psi},$ (44)
where $J\equiv\frac{\partial X^{\pm}}{\partial x^{\pm}}$ denotes (from now) the Jacobian of the coordinate transformation linking the coordinates $x^{\pm}$ and $X^{\pm}$. It is clear from the previous definition that $\braket{1_{\pm}2_{\pm}3_{\pm}4_{\pm}}^{x}_{\psi{\rm ren}}$ is a tensor transforming covariantly under changes of $x^{\pm}$.
Our task is to express $\braket{1_{\pm}2_{\pm}3_{\pm}4_{\pm}}^{x}_{\psi{\rm ren}}$ in terms of normal ordered quantities defined in an arbitrary representation so that we get an expression that is easy to compute with in some situations where a certain ‘vacuum state’ might capture a particular idealized physical situation. We then use (B.2) —written for the representation constructed in terms of the Minkowskian coordindates $X^{\pm}$—to write $\braket{:1_{\pm}2_{\pm}3_{\pm}4_{\pm}\!:}^{X}_{\psi}$ in terms of $\braket{1_{\pm}2_{\pm}3_{\pm}4_{\pm}}^{X}_{\psi}$, namely
$\braket{1_{\pm}2_{\pm}3_{\pm}4_{\pm}}^{x}_{\psi{\rm ren}}\equiv J_{1}J_{2}J_{3}J_{4}\braket{:1_{\pm}2_{\pm}3_{\pm}4_{\pm}\!:}^{X}_{\psi}=J_{1}J_{2}J_{3}J_{4}\braket{1_{\pm}2_{\pm}3_{\pm}4_{\pm}}^{X}_{\psi}{}$
$-J_{1}J_{2}J_{3}J_{4}\left(\braket{:1_{\pm}2_{\pm}\!:}^{X}_{\psi}\xi_{34}^{X^{\pm}}+\braket{:1_{\pm}3_{\pm}\!:}^{X}_{\psi}\xi_{24}^{X^{\pm}}+\braket{:1_{\pm}4_{\pm}\!:}^{X}_{\psi}\xi_{23}^{X^{\pm}}+\right$.{}
$\left.+\braket{:2_{\pm}3_{\pm}\!:}^{X}_{\psi}\xi_{14}^{X^{\pm}}+\braket{:2_{\pm}4_{\pm}\!:}^{X}_{\psi}\xi_{13}^{X^{\pm}}+\braket{:3_{\pm}4_{\pm}\!:}^{X}_{\psi}\xi_{12}^{X^{\pm}}\right){}$
$-J_{1}J_{2}J_{3}J_{4}\left(\xi_{12}^{X^{\pm}}\xi_{34}^{X^{\pm}}+\xi_{13}^{X^{\pm}}\xi_{24}^{X^{\pm}}+\xi_{14}^{X^{\pm}}\xi_{23}^{X^{\pm}}\right),$ (45)
Now equation (43) implies
$\braket{1_{\pm}2_{\pm}3_{\pm}4_{\pm}}^{x}_{\psi{\rm ren}}\equiv J_{1}J_{2}J_{3}J_{4}\braket{:1_{\pm}2_{\pm}3_{\pm}4_{\pm}\!:}^{X}_{\psi}=\braket{1_{\pm}2_{\pm}3_{\pm}4_{\pm}}^{x}_{\psi}{}$
$-J_{1}J_{2}J_{3}J_{4}\left(\braket{:1_{\pm}2_{\pm}\!:}^{X}_{\psi}\xi_{34}^{X^{\pm}}+\braket{:1_{\pm}3_{\pm}\!:}^{X}_{\psi}\xi_{24}^{X^{\pm}}+\braket{:1_{\pm}4_{\pm}\!:}^{X}_{\psi}\xi_{23}^{X^{\pm}}+\right$.{}
$\left.+\braket{:2_{\pm}3_{\pm}\!:}^{X}_{\psi}\xi_{14}^{X^{\pm}}+\braket{:2_{\pm}4_{\pm}\!:}^{X}_{\psi}\xi_{13}^{X^{\pm}}+\braket{:3_{\pm}4_{\pm}\!:}^{X}_{\psi}\xi_{12}^{X^{\pm}}\right){}$
$-J_{1}J_{2}J_{3}J_{4}\left(\xi_{12}^{X^{\pm}}\xi_{34}^{X^{\pm}}+\xi_{13}^{X^{\pm}}\xi_{24}^{X^{\pm}}+\xi_{14}^{X^{\pm}}\xi_{23}^{X^{\pm}}\right),$ (46)
where we used (B.2)—written for the representation constructed in terms of the Minkowskian coordindates $X^{\pm}$—to write $\braket{:1_{\pm}2_{\pm}3_{\pm}4_{\pm}\!:}^{X}_{\psi}$ in terms of $\braket{1_{\pm}2_{\pm}3_{\pm}4_{\pm}}^{X}_{\psi}$ followed by equation (43). We can write the previous equation as follows
$\braket{1_{\pm}2_{\pm}3_{\pm}4_{\pm}}^{x}_{\psi{\rm ren}}=\braket{1_{\pm}2_{\pm}3_{\pm}4_{\pm}}^{x}_{\psi}{}$
$-J_{1}J_{2}J_{3}J_{4}\left[(\braket{1_{\pm}2_{\pm}}^{X}_{\psi}-\xi_{12}^{X^{\pm}})\xi_{34}^{X^{\pm}}+(\braket{1_{\pm}3_{\pm}}^{X}_{\psi}-\xi_{13}^{X^{\pm}})\xi_{24}^{X^{\pm}}+(\braket{1_{\pm}4_{\pm}}^{X}_{\psi}-\xi_{14}^{X^{\pm}})\xi_{23}^{X^{\pm}}\right$.{}
$\left.+(\braket{2_{\pm}3_{\pm}}^{X}_{\psi}-\xi_{23}^{X^{\pm}})\xi_{14}^{X^{\pm}}+(\braket{2_{\pm}4_{\pm}}^{X}_{\psi}-\xi_{24}^{X^{\pm}})\xi_{13}^{X^{\pm}}+(\braket{3_{\pm}4_{\pm}}^{X}_{\psi}-\xi_{34}^{X^{\pm}})\xi_{12}^{X^{\pm}}\right]{}$
$-J_{1}J_{2}J_{3}J_{4}\left(\xi_{12}^{X^{\pm}}\xi_{34}^{X^{\pm}}+\xi_{13}^{X^{\pm}}\xi_{34}^{X^{\pm}}+\xi_{14}^{X^{\pm}}\xi_{23}^{X^{\pm}}\right),$ (47)
using that for instance $J_{1}J_{2}\braket{1_{\pm}2_{\pm}}^{X}_{\psi}=\braket{1_{\pm}2_{\pm}}^{x}_{\psi}$ and denoting terms like $J_{1}J_{2}\xi_{12}^{X^{\pm}}=J(\xi_{12}^{X^{\pm}})$ we obtain
$\braket{1_{\pm}2_{\pm}3_{\pm}4_{\pm}}^{x}_{\psi{\rm ren}}=\braket{1_{\pm}2_{\pm}3_{\pm}4_{\pm}}^{x}_{\psi}{}$
$-\left[(-J(\xi_{12}^{X^{\pm}}))J(\xi_{34}^{X^{\pm}})+(\braket{1_{\pm}3_{\pm}}^{x}_{\psi}-J(\xi_{13}^{X^{\pm}}))J(\xi_{24}^{X^{\pm}})+(\braket{1_{\pm}4_{\pm}}^{x}_{\psi}-J(\xi_{14}^{X^{\pm}}))J(\xi_{23}^{X^{\pm}})\right$.{}
$\left.+(\braket{2_{\pm}3_{\pm}}^{x}_{\psi}-J(\xi_{23}^{X^{\pm}}))J(\xi_{14}^{X^{\pm}})+(\braket{2_{\pm}4_{\pm}}^{x}_{\psi}-J(\xi_{24}^{X^{\pm}}))J(\xi_{13}^{X^{\pm}})+(\braket{3_{\pm}4_{\pm}}^{x}_{\psi}-J(\xi_{34}^{X^{\pm}}))J(\xi_{12}^{X^{\pm}})\right]{}$
$-J(\xi_{12}^{X^{\pm}})J(\xi_{34}^{X^{\pm}})-J(\xi_{13}^{X^{\pm}})J(\xi_{24}^{X^{\pm}})-J(\xi_{14}^{X^{\pm}})J(\xi_{23}^{X^{\pm}}){}$
$=\braket{1_{\pm}2_{\pm}3_{\pm}4_{\pm}}^{x}_{\psi}{}$
$-\braket{1_{\pm}2_{\pm}}^{x}_{\psi}J(\xi_{34}^{X^{\pm}})-\braket{1_{\pm}3_{\pm}}^{x}_{\psi}J(\xi_{24}^{X^{\pm}})-\braket{1_{\pm}4_{\pm}}^{x}_{\psi}J(\xi_{23}^{X^{\pm}}){}$
$-\braket{2_{\pm}3_{\pm}}^{x}_{\psi}J(\xi_{14}^{X^{\pm}})-\braket{2_{\pm}4_{\pm}}^{x}_{\psi}J(\xi_{13}^{X^{\pm}})-\braket{3_{\pm}4_{\pm}}^{x}_{\psi}J(\xi_{12}^{X^{\pm}}){}$
$+J(\xi_{12}^{X^{\pm}})J(\xi_{34}^{X^{\pm}})+J(\xi_{13}^{X^{\pm}})J(\xi_{24}^{X^{\pm}})+J(\xi_{14}^{X^{\pm}})J(\xi_{23}^{X^{\pm}})+J(\xi_{23}^{X^{\pm}})J(\xi_{14}^{X^{\pm}})+J(\xi_{24}^{X^{\pm}})J(\xi_{13}^{X^{\pm}})+J(\xi_{34}^{X^{\pm}})J(\xi_{12}^{X^{\pm}}){}$
$-J(\xi_{12}^{X^{\pm}})J(\xi_{34}^{X^{\pm}})-J(\xi_{13}^{X^{\pm}})J(\xi_{24}^{X^{\pm}})-J(\xi_{14}^{X^{\pm}})J(\xi_{23}^{X^{\pm}}){}$
$=\braket{1_{\pm}2_{\pm}3_{\pm}4_{\pm}}^{x}_{\psi}{}$
$-\braket{1_{\pm}2_{\pm}}^{x}_{\psi}J(\xi_{34}^{X^{\pm}})-\braket{1_{\pm}3_{\pm}}^{x}_{\psi}J(\xi_{24}^{X^{\pm}})-\braket{1_{\pm}4_{\pm}}^{x}_{\psi}J(\xi_{23}^{X^{\pm}}){}$
$-\braket{2_{\pm}3_{\pm}}^{x}_{\psi}J(\xi_{14}^{X^{\pm}})-\braket{2_{\pm}4_{\pm}}^{x}_{\psi}J(\xi_{13}^{X^{\pm}})-\braket{3_{\pm}4_{\pm}}^{x}_{\psi}J(\xi_{12}^{X^{\pm}}){}$
$+J(\xi_{12}^{X^{\pm}})J(\xi_{34}^{X^{\pm}})+J(\xi_{13}^{X^{\pm}})J(\xi_{24}^{X^{\pm}})+J(\xi_{14}^{X^{\pm}})J(\xi_{23}^{X^{\pm}})$ (48)
[…]
$\braket{1_{\pm}2_{\pm}3_{\pm}4_{\pm}}^{x}_{\psi{\rm ren}}=\braket{:1_{\pm}2_{\pm}3_{\pm}4_{\pm}\!:}^{x}_{\psi}$ (49)
$+\braket{:1_{\pm}2_{\pm}\!:}^{x}_{\psi}\xi_{34}^{x^{\pm}}+\braket{:1_{\pm}3_{\pm}\!:}^{x}_{\psi}\xi_{24}^{x^{\pm}}+\braket{:1_{\pm}4_{\pm}\!:}^{x}_{\psi}\xi_{23}^{x^{\pm}}+\braket{:2_{\pm}3_{\pm}\!:}^{x}_{\psi}\xi_{14}^{x^{\pm}}+\braket{:2_{\pm}4_{\pm}\!:}^{x}_{\psi}\xi_{13}^{x}+\braket{:3_{\pm}4_{\pm}\!:}^{x}_{\psi}\xi_{12}^{x^{\pm}}+{}$
$+\xi_{12}^{x^{\pm}}\xi_{34}^{x^{\pm}}+\xi_{13}^{x}\xi_{24}^{x^{\pm}}+\xi_{14}^{x^{\pm}}\xi_{23}^{x^{\pm}}{}$
$-(\braket{:1_{\pm}2_{\pm}\!:}^{x}_{\psi}+\xi_{12}^{x^{\pm}})J(\xi_{34}^{X^{\pm}})-(\braket{:1_{\pm}3_{\pm}\!:}^{x}_{\psi}+\xi_{13}^{x})J(\xi_{24}^{X^{\pm}})-(\braket{:1_{\pm}4_{\pm}\!:}^{x}_{\psi}+\xi_{14}^{x^{\pm}})J(\xi_{23}^{X^{\pm}}){}$
$-(\braket{:2_{\pm}3_{\pm}\!:}^{x}_{\psi}+\xi_{23}^{x^{\pm}})J(\xi_{14}^{X^{\pm}})-(\braket{:2_{\pm}4_{\pm}\!:}^{x}_{\psi}+\xi_{24}^{x^{\pm}})J(\xi_{13}^{X^{\pm}})-(\braket{:3_{\pm}4_{\pm}\!:}^{x}_{\psi}+\xi_{34}^{x^{\pm}})J(\xi_{12}^{X^{\pm}}){}$
$+J(\xi_{12}^{X^{\pm}})J(\xi_{34}^{X^{\pm}})+J(\xi_{13}^{X^{\pm}})J(\xi_{24}^{X^{\pm}})+J(\xi_{14}^{X^{\pm}})J(\xi_{23}^{X^{\pm}}){}$
$=\braket{:1_{\pm}2_{\pm}3_{\pm}4_{\pm}\!:}^{x}_{\psi}{}$
$+\braket{:1_{\pm}2_{\pm}\!:}^{x}_{\psi}\left[\xi_{34}^{x^{\pm}}-J(\xi_{34}^{X^{\pm}})\right]+\braket{:1_{\pm}3_{\pm}\!:}^{x}_{\psi}[\xi_{24}^{x^{\pm}}-J(\xi_{24}^{X^{\pm}})]+\braket{:1_{\pm}4_{\pm}\!:}^{x}_{\psi}[\xi_{23}^{x^{\pm}}-J(\xi_{23}^{X^{\pm}})]{}$
$+\braket{:2_{\pm}3_{\pm}\!:}^{x}_{\psi}[\xi_{14}^{x^{\pm}}-J(\xi_{14}^{X^{\pm}})]+\braket{:2_{\pm}4_{\pm}\!:}^{x}_{\psi}[\xi_{13}^{x^{\pm}}-J(\xi_{13}^{X^{\pm}})]+\braket{:3_{\pm}4_{\pm}\!:}^{x}_{\psi}[\xi_{12}^{x^{\pm}}-J(\xi_{12}^{X^{\pm}})]+{}$
$+\xi_{12}^{x^{\pm}}\xi_{34}^{x^{\pm}}+\xi_{13}^{x}\xi_{24}^{x^{\pm}}+\xi_{14}^{x^{\pm}}\xi_{23}^{x^{\pm}}{}$
$-\xi_{12}^{x^{\pm}}J(\xi_{34}^{X^{\pm}})-\xi_{13}^{x}J(\xi_{24}^{X^{\pm}})-\xi_{14}^{x^{\pm}}J(\xi_{23}^{X^{\pm}})-\xi_{23}^{x^{\pm}}J(\xi_{14}^{X^{\pm}})-\xi_{24}^{x^{\pm}}J(\xi_{13}^{X^{\pm}})-\xi_{34}^{x^{\pm}}J(\xi_{12}^{X^{\pm}}){}$
$+J(\xi_{12}^{X^{\pm}})J(\xi_{34}^{X^{\pm}})+J(\xi_{13}^{X^{\pm}})J(\xi_{24}^{X^{\pm}})+J(\xi_{14}^{X^{\pm}})J(\xi_{23}^{X^{\pm}}){}$
$=\braket{:1_{\pm}2_{\pm}3_{\pm}4_{\pm}\!:}^{x}_{\psi}+\braket{:1_{\pm}2_{\pm}\!:}^{x}_{\psi}\left[\xi_{34}^{x^{\pm}}-J(\xi_{34}^{X^{\pm}})\right]+\braket{:1_{\pm}3_{\pm}\!:}^{x}_{\psi}[\xi_{24}^{x^{\pm}}-J(\xi_{24}^{X^{\pm}})]+\braket{:1_{\pm}4_{\pm}\!:}^{x}_{\psi}[\xi_{23}^{x^{\pm}}-J(\xi_{23}^{X^{\pm}})]+{}$
$+\braket{:2_{\pm}3_{\pm}\!:}^{x}_{\psi}[\xi_{14}^{x^{\pm}}-J(\xi_{14}^{X^{\pm}})]+\braket{:2_{\pm}4_{\pm}\!:}^{x}_{\psi}[\xi_{13}^{x}-J(\xi_{13}^{X^{\pm}})]+\braket{:3_{\pm}4_{\pm}\!:}^{x}_{\psi}[\xi_{12}^{x^{\pm}}-J(\xi_{12}^{X^{\pm}})]+[\xi_{12}^{x^{\pm}}-J(\xi_{12}^{X^{\pm}})][\xi_{34}^{x^{\pm}}-J(\xi_{34}^{X^{\pm}})]+[\xi_{13}^{x}-J(\xi_{13}^{X^{\pm}})][\xi_{24}^{x^{\pm}}-J(\xi_{24}^{X^{\pm}})]+[\xi_{14}^{x^{\pm}}-J(\xi_{14}^{X^{\pm}})][\xi_{23}^{x^{\pm}}-J(\xi_{23}^{X^{\pm}})]{}$
Writing the last line separately
$\braket{1_{\pm}2_{\pm}3_{\pm}4_{\pm}}^{x}_{\psi{\rm ren}}=\braket{:1_{\pm}2_{\pm}3_{\pm}4_{\pm}\!:}^{x}_{\psi}+\braket{:1_{\pm}2_{\pm}\!:}^{x}_{\psi}[\xi_{34}^{x^{\pm}}-J(\xi_{34}^{X^{\pm}})]+\braket{:1_{\pm}3_{\pm}\!:}^{x}_{\psi}[\xi_{24}^{x^{\pm}}-J(\xi_{24}^{X^{\pm}})]$ (50)
${}+\braket{:1_{\pm}4_{\pm}\!:}^{x}_{\psi}[\xi_{23}^{x^{\pm}}-J(\xi_{23}^{X^{\pm}})]+\braket{:2_{\pm}3_{\pm}\!:}^{x}_{\psi}[\xi_{14}^{x^{\pm}}-J(\xi_{14}^{X^{\pm}})]{}$
$+\braket{:2_{\pm}4_{\pm}\!:}^{x}_{\psi}[\xi_{13}^{x}-J(\xi_{13}^{X^{\pm}})]+\braket{:3_{\pm}4_{\pm}\!:}^{x}_{\psi}[\xi_{12}^{x^{\pm}}-J(\xi_{12}^{X^{\pm}})]+{}$
$+[\xi_{12}^{x^{\pm}}-J(\xi_{12}^{X^{\pm}})][\xi_{34}^{x^{\pm}}-J(\xi_{34}^{X^{\pm}})]+[\xi_{13}^{x}-J(\xi_{13}^{X^{\pm}})][\xi_{24}^{x^{\pm}}-J(\xi_{24}^{X^{\pm}})]+[\xi_{14}^{x^{\pm}}-J(\xi_{14}^{X^{\pm}})][\xi_{23}^{x^{\pm}}-J(\xi_{23}^{X^{\pm}})]{}$
Using that the quantities in the square brakets all have a well defined coincidence limit given by
$\lim_{1\to 2}\,[\xi_{12}^{x^{\pm}}-J(\xi_{12}^{X^{\pm}})]=-\frac{\hbar}{24\pi}\{X^{\pm},x^{\pm}\}.$ (51)
Note that all quanties are well defined in the coincidence limit so that by taking it we get the desired result:
$\braket{\psi|\partial_{\pm}\phi(x)\partial_{\pm}\phi(x)\partial_{\pm}\phi(x)\partial_{\pm}\phi(x)|\psi}_{\rm ren}={}$
$=\braket{\psi|:\partial_{\pm}\phi(x)\partial_{\pm}\phi(x)\partial_{\pm}\phi(x)\partial_{\pm}\phi(x):|\psi}-\frac{\hbar}{4\pi}\braket{\psi|:\partial_{\pm}\phi(x)\partial_{\pm}\phi(x):\psi}\{X,x\}+\frac{\hbar^{2}}{192\pi^{2}}\{X,x\}^{2}$ (52)
Finally, using (32) we get the following expression of the renormalized expectation value of the square of the energy momentum tensor Problema de signo creo en lo que sigue:
$\braket{\psi|T_{\pm\pm}T_{\pm\pm}|\psi}_{\rm ren}$ $\equiv$ $\braket{\psi|:T_{\pm\pm}T_{\pm\pm}:|\psi}$
- $\frac{\hbar}{2\pi}\braket{\psi|:T_{\pm\pm}:|\psi}\left(\partial_{\pm}\log(\Omega)\partial_{\pm}\log(\Omega)-\partial^{2}_{\pm}\log(\Omega)\right)+\frac{\hbar^{2}}{48\pi}\left(\partial_{\pm}\log(\Omega)\partial_{\pm}\log(\Omega)-\partial^{2}_{\pm}\log(\Omega)\right)^{2},{}$
whose covariance is granted by construction from the definition (44). The same method allows to write the $n$-point correlation functions for arbitrary $n$ in a close manner.
B.3 Calculation of $\braket{\psi|T_{++}T_{--}|\psi}_{\rm ren}$
In this case there are many terms that vanish in the Wick expansion of the relevant 4-point function due to the fact that $\braket{0_{x}|\partial_{+}\phi(x_{1})\partial_{-}\phi(x_{2})|0_{x}}=0$. We obtain
$\braket{1_{+}2_{+}3_{-}4_{-}}^{x}_{\psi}$ $=$ $\braket{:1_{+}2_{+}3_{-}4_{-}\!:}^{x}_{\psi}+\braket{:1_{+}2_{+}\!:}^{x}_{\psi}\xi_{34}^{x^{-}}+\braket{:3_{-}4_{-}\!:}^{x}_{\psi}\xi_{12}^{x^{+}}+\xi_{12}^{x^{+}}\xi_{34}^{x^{-}}.$ (54)
The renormalization prescription leads to
$\braket{1_{+}2_{+}3_{-}4_{-}}^{x}_{\psi{\rm ren}}\equiv J_{1}J_{2}J_{3}J_{4}\braket{:1_{+}2_{+}3_{-}4_{-}\!:}^{X}_{\psi}=$ (55)
${}=\braket{1_{+}2_{+}3_{-}4_{-}}^{x}_{\psi}-J_{1}J_{2}J_{3}J_{4}\left(\braket{:1_{+}2_{+}\!:}^{X}_{\psi}\xi_{34}^{X^{-}}+\braket{:3_{-}4_{-}\!:}^{X}_{\psi}\xi_{12}^{X^{+}}+\xi_{12}^{X^{+}}\xi_{34}^{X^{-}}\right)=$
$=\braket{1_{+}2_{+}3_{-}4_{-}}^{x}_{\psi}-J_{1}J_{2}J_{3}J_{4}\left(\braket{1_{+}2_{+}}^{X}_{\psi}\xi_{34}^{X^{-}}+\braket{3_{-}4_{-}}^{X}_{\psi}\xi_{12}^{X^{+}}+\xi_{12}^{X^{+}}\xi_{34}^{X^{-}}\right)+J_{1}J_{2}J_{3}J_{4}\left(\xi_{12}^{X^{+}}\xi_{34}^{X^{-}}+\xi_{34}^{X^{-}}\xi_{12}^{X^{+}}\right){}=$
${}=\braket{1_{+}2_{+}3_{-}4_{-}}^{x}_{\psi}-\braket{1_{+}2_{+}}^{x}_{\psi}J(\xi_{34}^{X^{-}})-\braket{3_{-}4_{-}}^{x}_{\psi}J(\xi_{12}^{X^{+}})+J(\xi_{12}^{X^{+}})J(\xi_{34}^{X^{-}})=$
$=\braket{:1_{+}2_{+}3_{-}4_{-}\!:}^{x}_{\psi}+\braket{:1_{+}2_{+}\!:}^{x}_{\psi}\xi_{34}^{x^{-}}+\braket{:3_{-}4_{-}\!:}^{x}_{\psi}\xi_{12}^{x^{+}}+\xi_{12}^{x^{+}}\xi_{34}^{x^{-}}{}$
$-\braket{:1_{+}2_{+}:}^{x}_{\psi}J(\xi_{34}^{X^{-}})-\braket{:3_{-}4_{-}:}^{x}_{\psi}J(\xi_{12}^{X^{+}})-\xi_{12}^{x^{+}}J(\xi_{34}^{X^{-}})-\xi_{34}^{x^{-}}J(\xi_{12}^{X^{+}})+J(\xi_{12}^{X^{+}})J(\xi_{34}^{X^{-}}){}=$
$=\braket{:1_{+}2_{+}3_{-}4_{-}\!:}^{x}_{\psi}+\braket{:1_{+}2_{+}\!:}^{x}_{\psi}(\xi_{34}^{x^{-}}-J(\xi_{34}^{X^{-}})+\braket{:3_{-}4_{-}\!:}^{x}_{\psi}(\xi_{12}^{x^{+}}-J(\xi_{12}^{X^{+}})+(\xi_{12}^{x^{+}}-J(\xi_{12}^{X^{+}})(\xi_{34}^{x^{-}}-J(\xi_{34}^{X^{-}}),{}$
where we have followed step by step what we did previously. Thus,
$\braket{\psi|\partial_{+}\phi(x)\partial_{+}\phi(x)\partial_{-}\phi(x)\partial_{-}\phi(x)|\psi}_{\rm ren}={}$
$=\braket{\psi|:\partial_{+}\phi(x)\partial_{+}\phi(x)\partial_{-}\phi(x)\partial_{-}\phi(x):|\psi}-\frac{\hbar}{24\pi}\braket{\psi|:\partial_{+}\phi(x)\partial_{+}\phi(x):\psi}\{X^{-},x^{-}\}{}$
$-\frac{\hbar}{24\pi}\braket{\psi|:\partial_{-}\phi(x)\partial_{-}\phi(x):\psi}\{X^{+},x^{+}\}+\frac{\hbar^{2}}{576\pi^{2}}\{X^{+},x^{+}\}\{X^{-},x^{-}\}.$ (56)
And finally, using (32), we get
$\braket{\psi|T_{++}T_{--}|\psi}_{\rm ren}=\braket{\psi|:T_{++}T_{--}:|\psi}{}$
$+\frac{\hbar}{12\pi}\braket{\psi|:T_{++}:\psi}\left(\partial_{-}\log(\Omega)\partial_{-}\log(\Omega)-\partial^{2}_{-}\log(\Omega)\right)+\frac{\hbar}{12\pi}\braket{\psi|:T_{--}:\psi}\left(\partial_{+}\log(\Omega)\partial_{+}\log(\Omega)-\partial^{2}_{+}\log(\Omega)\right){}$
$+\frac{\hbar^{2}}{144\pi^{2}}\left(\partial_{-}\log(\Omega)\partial_{-}\log(\Omega)-\partial^{2}_{-}\log(\Omega)\right)\left(\partial_{+}\log(\Omega)\partial_{+}\log(\Omega)-\partial^{2}_{+}\log(\Omega)\right).$ (57)
Applying the previous equation to the vacuum states associated with the notion of normal ordering involved,
$\braket{0|T_{++}T_{--}|0}_{\rm ren}=\frac{\hbar^{2}}{144\pi^{2}}\left(\partial_{-}\log(\Omega)\partial_{-}\log(\Omega)-\partial^{2}_{-}\log(\Omega)\right)\left(\partial_{+}\log(\Omega)\partial_{+}\log(\Omega)-\partial^{2}_{+}\log(\Omega)\right){}$
$=\braket{0|T_{++}|0}_{\rm ren}\braket{0|T_{--}|0}_{\rm ren}$ (58)
We see that the large ‘vacuum fluctuations’ are not going to manifest in such correlation functions. The reason is that $T_{++},T_{--}]=0$ and the conformal vacuum states are product states.
B.4 Operator product expansion
One can express the previous results in the language of the OPE. For the energy momentum tensor, the relevant expansion is
$\braket{\psi|\partial_{x^{\pm}}\phi_{1}\partial_{x^{\pm}}\phi_{2}|\psi}$ $=$ $J_{1}J_{2}\braket{\psi|\partial_{X^{\pm}}\phi_{1}\partial_{X^{\pm}}\phi_{2}|\psi}{}$ (59)
$\approx$ $J_{1}J_{2}\left(\braket{\psi|:\partial_{X^{\pm}}\phi_{1}\partial_{X^{\pm}}\phi_{2}\!:_{X}|\psi}\underbrace{-\frac{\hbar}{12\pi}\partial_{X^{\pm}_{1}}\partial_{X^{\pm}_{2}}\log\left({|(X^{+}_{1}-X^{+}_{2})(X^{-}_{1}-X^{-}_{2})|}\right)}_{\braket{0_{X}|\partial_{\pm}\phi(X_{1})\partial_{\pm}\phi(X_{2})|0_{X}}=-\frac{\hbar}{12\pi}\frac{1}{|X^{\pm}_{1}-X^{\pm}_{2}|^{2}}}\right){}$
$\approx$ $\braket{\psi|(\partial_{x^{\pm}}\phi)^{2}|\psi}_{\rm ren}-\frac{\hbar}{12\pi}\partial_{x_{1}^{\pm}}\partial_{x_{2}^{\pm}}\log\left({\sigma(x_{1},x_{2})},\right)$
where $\sigma(x_{1},x_{2})$ is the geodesic distance squared between $x_{1}$ and $x_{2}$, and $\approx$ denotes the equality when $x_{1}\to x_{2}$. Similarly
$\braket{\psi|\partial_{x^{\pm}}\phi_{1}\partial_{x^{\pm}}\phi_{2}\partial_{x^{\pm}}\phi_{3}\partial_{x^{\pm}}\phi_{4}|\psi}$ $\approx$ $\braket{\psi|(\partial_{x^{\pm}}\phi)^{4}|\psi}_{\rm ren}$
- ${}\frac{\hbar}{2\pi}\braket{\psi|(\partial_{x^{\pm}}\phi)^{2}|\psi}_{\rm ren}\partial_{x_{1}^{\pm}}\partial_{x_{2}^{\pm}}\log\left({\sigma(x_{1},x_{2})}\right)$
$+$ ${}\frac{\hbar^{2}}{48\pi}\partial_{x_{1}^{\pm}}\partial_{x_{2}^{\pm}}\partial_{x_{3}^{\pm}}\partial_{x_{4}^{\pm}}\log\left({\sigma(x_{1},x_{2})}\right)\log\left({\sigma(x_{3},x_{4})}\right).$
One can express the OPE for any correlation function of primary fields along the same lines.
Appendix C Relation to 4d
The Klein-Gordon equation $\square\Phi-m^{2}\Phi=0$ in the Schwarzschild background in spherical coordinates can be written as
$4\partial_{u}\partial_{v}\phi_{\ell}+\underbrace{\left(1-\frac{2M}{r}\right)\left(\frac{\ell(\ell+1)}{r^{2}}+\frac{2M}{r^{3}}+m^{2}\right)}_{V_{\ell}(r)}\phi_{\ell}=0,$ (61)
where we used ansatz $\Phi(u,v,\theta,\varphi)=\phi_{\ell}(u,v)Y_{\ell m}(\theta,\varphi)/r$ with $u=t-r_{*}$, $v=t+r_{*}$, $r_{*}=r+2M\log(r/(2M)-1)$ and the radial coordinate should be taken as an implicit function of the null coordinates, $r=r(u,v)$. At the horizon $r=2M$ the previous equation becomes the 2d conformally invariant field equation of a massless scalar field for all the modes, namely
$\partial_{u}\partial_{v}\phi_{\ell}=0.$ (62)
The non locality of quantum field theory implies that the effects of the potential $V_{\ell}(r)$ cannot be neglected in discussion involving the Hilbert space construction and the definition of some particular states of interest. However, qualitative information about the quantum theory can be obtained especially if we simplify the discussion further by focusing on the spherically symmetric modes $\ell=0$ (which is the main channel for scalar Hawking radiation Page:1976df ). Identifying $\phi_{0}$ with the 2d scalar field $\phi$ of Section I we have that the contribution of these modes to the relevant components of the energy momentum tensor of the Klein-Gordon field is
$\left.T^{\rm sph}_{vv}\right|_{\rm H}=\left.\partial_{v}\Phi\partial_{v}\Phi\right|_{\rm H}=\left.\left(\frac{\partial_{v}\phi}{r}-\frac{\phi}{r^{2}}\frac{\partial r}{\partial v}\right)\left(\frac{\partial_{v}\phi}{r}-\frac{\phi}{r^{2}}\frac{\partial r}{\partial v}\right)\right|_{\rm H}=\left.\frac{\partial_{v}\phi\partial_{v}\phi}{4M^{2}}\right|_{\rm H}=\frac{T^{\rm 2d}_{vv}}{4M^{2}}.$ (63)